OpenStudy (anonymous):

Simplify 3√242x^5y^5) + 11xy√2x^3y^3)

OpenStudy (anonymous):

\[44\sqrt{2x ^{5}y ^{5}}\]

OpenStudy (anonymous):

but how do you simplify it like that?

OpenStudy (anonymous):

1. \[33\sqrt{2}\sqrt{x ^{5}y ^{5}}+11xy \sqrt{2} x ^{1.5}y ^{1.5}=33\sqrt{2x ^{5}y ^{5}}+11\sqrt{2x^5}\]

OpenStudy (anonymous):

\[(33+11)\sqrt{2x ^{5}y ^{5}}=44\sqrt{2}x ^{5}y ^{5}\]

OpenStudy (anonymous):

Any question?

OpenStudy (anonymous):

Mr. Mike is everything ok? ... Have I missed sth. or misunderstand sth in the question. thanks.

OpenStudy (anonymous):

no thanks i was eating sorry